OpenStudy (anonymous):

Write the equation of the circle containing the points J(-6, 0), K(-3, 3), and L(0, 0). Show all work to receive credit.

OpenStudy (anonymous):

The slope of JK is 1. The center of the circle is (-3,0) Which means the radius is the distance between (-3,0) and (-6,0) D=sqrt of (-6(-3))^2+(0-0)^2 =sqrt of -3^2 =sqrt of 9 r=3

OpenStudy (anonymous):

What am I missing?

OpenStudy (anonymous):

the equation of a circle with center at (x1,y1) is (x-x1)^2 + (y-y1)^2 = r^2 you have three points and three unknowns find x1,y1 and r

OpenStudy (anonymous):

the equation of a circle with radius 3 and center at (-3,0) is (x+3)^2+y^2 = 9

OpenStudy (anonymous):

okay so what do i do after (x-3)^2+y^2=9? i mean how do i solve that?

OpenStudy (anonymous):

you are asked to write the equation of the circle. That IS the equation of the circle. You don't have to solve for anything

OpenStudy (anonymous):

its like asking to write the equation of a line passing through the origin and slope 2 y =2x is the equation of that line. you don't have to solve that equation.

OpenStudy (anonymous):

btw, the equation is (x+3)^2+y^2 = 9, not (x-3)^2+y^2=9

OpenStudy (anonymous):

Oh okay so i just write it?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

you are welcome